What is the the frequency (in kHz) of radio wave radiation that has a wavelength of 6.870e+5 micrometers? The speed of light is 3.00x10 8 m/s.
1
Expert's answer
2020-12-21T03:30:52-0500

λ = C/f = 3.00x10 8 m/s / 0,687 m = 436 681 223 Hz. = 436 681, 223 kHz